# Feed Two Birds with One Scone: Exploiting Wild Data for Both Out-of-Distribution Generalization and Detection

### v1

@@ 4.3 Qualitative Insights
Figure 3 shows t-SNE visualizations (Van der Maaten & Hinton, 2008) of the penultimate-layer feature embedding, for WOODS (c) vs. our method (d). The embeddings are extracted from the test split. The blue points denote the test ID set (CIFAR-10), the green points are the test samples from CIFAR-10-C, and the gray points are from SVHN. This visualization suggests that embeddings of CIFAR and CIFAR-C become more aligned with our margin-based optimization, which arguably leads to improved OOD generalization performance. This observation corroborates our theoretical insight in Section 3.3.
@@ 4.4 Experiments on ImageNet
In this section, we provide additional large-scale results on the ImageNet benchmark. We use ImageNet-100 as the in-distribution data ($\mathbb{P}_{\text{in}}$), with labels provided in Appendix C. For the covariate-shifted OOD data ($\mathbb{P}_{\text{out}}^{\text{covariate}}$), we use ImageNet-100-C with Gaussian noise in the experiment. For the semantic-shifted OOD data, we use the high-resolution natural images from iNaturalist (Van Horn et al., 2018), with the same subset as MOS (Huang & Li, 2021). The wild data $\mathbb{P}_{\text{in}}$ is a mixture of ID data, covariate-shifted data ($\pi_{c}=0.5$), and semantic shifted data ($\pi_{s}=0.1$). We fine-tune ResNet-34 (He et al., 2016) (pre-trained on ImageNet) for 100 epochs, with an initial learning rate of 0.01 and a batch size of 64. Results in Table 5 suggest that our method can improve both ID and OOD accuracy compared to WOODS (the most competitive baseline).
@@ 5 Related Works
is of vital importance for machine learning models deployed in the open world. Recent advances in OOD detection can be broadly categorized into post hoc and regularization-based methods. In particular, post hoc methods (Hendrycks & Gimpel, 2017; Liang et al., 2018; Lee et al., 2018; Liu et al., 2020; Huang et al., 2021; Sun et al., 2021, 2022) focus on deriving test-time OOD scoring functions for a pre-trained classifier. Our proposed work is closer to another line of work (Bevandić et al., 2018; Hendrycks et al., 2018; Malinin & Gales, 2018; Liu et al., 2020; Du et al., 2021; Ming et al., 2022), which addresses the OOD detection problem by training-time regularization. For example, models are encouraged to give predictions with lower confidence (Hendrycks et al., 2018) or higher energies (Liu et al., 2020). These methods require access to a clean OOD dataset for training, which can be restrictive. To circumvent this, a recent work WOODS (Katz-Samuels et al., 2022) first explored using wild mixture data consisting of the ID and semantically shifted OOD data $\mathbb{P}_{\text{wild}}:=(1-\pi)\mathbb{P}_{\text{in}}+\pi\mathbb{P}_{\text{out}}^{\text{semantic}}$. The crucial difference between our work and WOODS is whether the wild mixture data contains covariate-shifted data, which introduces new challenges not considered in prior work. As we show in this work, our formulation uniquely enables both OOD generalization and OOD detection, in one coherent framework.
@@
is a fundamental problem in machine learning, which aims to generalize to covariate-shifted data without any sample from the target domain (Muandet et al., 2013; Arjovsky et al., 2019; Bahng et al., 2020; Wang et al., 2022b; Xie et al., 2021). OOD generalization is more challenging compared to the classic domain adaptation problem (Daume III & Marcu, 2006; Blitzer et al., 2007; Ben-David et al., 2010; Ganin & Lempitsky, 2015; Tzeng et al., 2017; Redko et al., 2019; Kang et al., 2019; Kumar et al., 2020; Wang et al., 2022c), which assumes access to labeled samples from the target domain.
To highlight a few works, IRM (Arjovsky et al., 2019) and its variants (Krueger et al., 2021; Ahuja et al., 2020) aim to find invariant representation from different training environments via an invariant risk regularizer. GroupDRO (Sagawa et al., 2019) and Probabilistic Group DRO (Ghosal & Li, 2023) minimize the worst-case training loss over a set of groups. Zhou et al. (2020) propose generating outlier samples of a novel domain, which are then used for improving the generalization of the classifier. Besides algorithm innovation, benchmark efforts have also been pursued by DomainBed (Gulrajani & Lopez-Paz, 2020) and OoD-Bench (Ye et al., 2022), which facilitates evaluation on OOD generalization. OS-SDG (Zhu & Li, ) relies on one source domain with labels to train the model, while our framework can exploit unlabeled wild data naturally arising in the wild, which is a mixed composition of three data distributions. Different from previous literature, we focus on improving OoD robustness in classifiers by learning from the wild mixture data and building an OOD detector at the same time. To the best of our knowledge, we are the first work that leverages wild data for both OOD generalization and OOD detection purposes. Our framework also uniquely allows leveraging covariate-shifted data freely arising in the wild, without requiring any labeling.
@@
aims to leverage labeled data from a related domain (source domain) and improve the model performance for the target domain, where there exists category gap for label sets between the source and target domains (You et al., 2019). Several works have been proposed to address this problem (Saito et al., 2020; Fu et al., 2020; Li et al., 2021; Chen et al., 2022; Kundu et al., 2022; Chang et al., 2022; Garg et al., 2022). UAN (You et al., 2019) presents a universal adaptation network that exploits both the domain similarity and prediction uncertainty of each sample for promoting common-class adaptation. DANCE (Saito et al., 2020) proposes a domain adaptative neighborhood clustering technique for category shift-agnostic adaptation via entropy optimization. The work in Chang et al. (2022) proposes a unified optimal transport-based framework to encourage both global cluster discrimination and local consistency of samples. Different from prior works, we leverage both labeled in-distribution data and unlabeled wild data when training our model. Such unlabeled wild data naturally arise in real-world environments and have not been considered in prior literature.
@@
is a classic machine learning problem, which aims to learn classifiers from positive and unlabeled data (Letouzey et al., 2000). Multiple prior works have been proposed for discussing PU learning (Hsieh et al., 2015; Zhao et al., 2022; Acharya et al., 2022; Chapel et al., 2020; Xu & Denil, 2021). The work in (Niu et al., 2016) proposes a theoretical comparison of positive-unlabeled learning against positive-negative learning based on the upper bounds of estimation errors.  Du Plessis et al. (2015) presents a convex formulation for PU learning by using different loss functions for positive and unlabeled samples. Margin-based PU learning (Gong et al., 2018) introduces a provable positive margin-based PU learning algorithm for classification under the truncated linear distributions. There are two key differences between ours and PU learning: (1) PU learning only considers the task of distinguishing $\mathbb{P}_{\text{out}}$ (anomalous) and $\mathbb{P}_{\text{in}}$ (normal), not the task of doing classification simultaneously. We consider OOD detection which additionally requires learning a classifier for the distribution $\mathbb{P}_{\mathcal{X}\mathcal{Y}}$. (2) PU learning does not consider the generalization aspect under covariate-shifted OOD data, whereas our framework handles it in addition to semantic-shifted OOD data.
@@ Appendix B Experimental Details
margin OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$ $\tilde{\text{out}}$%
@@
We use Wide ResNet (Zagoruyko & Komodakis, 2016) with 40 layers and widen factor of 2. For evaluating the post hoc OOD detection baselines, we use the model trained with the CE loss on $\mathbb{P}_{\text{in}}$. We employ the same pre-trained model from the github: https://github.com/wetliu/energy_ood, which was trained on the complete CIFAR-10 dataset. Specifically, the model is trained using cross-entropy loss for 200 epochs. The learning rate is started at 0.1 and then decays by multiplier 0.1 at 100, 150, and 175 epochs. To facilitate easy comparison, the pre-trained model and baseline results are consistent with (Katz-Samuels et al., 2022) (courtesy of Table 4).
@@
For OOD generalization baselines, we use the same network architecture, Wide ResNet-40-2 (Zagoruyko & Komodakis, 2016), and train from scratch using respective losses. The baselines in Table 3 are trained on CIFAR-10 (Krizhevsky et al., 2009) $50,000$ labeled training examples. We follow the default hyperparameter setting as in original papers whenever applicable. For Mixup (Zhang et al., 2018), we follow the original paper and set hyperparameter $\alpha=1$ to control the strength of interpolation between feature-target pairs. The $\lambda$ for IRM (Arjovsky et al., 2019) baseline is set to 100 and the $\lambda$ for VREx (Krueger et al., 2021) is set to 10 for the penalty weights. Following the original training configuration in WRN (Zagoruyko & Komodakis, 2016), all the baselines are trained for 200 epochs with batch size 128. We use the SGD optimizer with an initial learning rate of 0.1. The learning rate decays by a factor of 10 after 60, 120, and 180 epochs. Weight decay is set to $10^{-4}$. All models are implemented in PyTorch 1.8.1. We evaluate the trained model on the CIFAR-10 test set (ID accuracy) and CIFAR-10-C (OOD accuracy).
@@ Appendix C Details of Datasets
MNIST (LeCun, 1998) is a large database of handwritten digits with 10 categories and is widely used in the field of machine learning. The MNIST contains $60,000$ training images and $10,000$ test images.
@@
CIFAR-10 (Krizhevsky et al., 2009) contains $60,000$ color images with 10 classes. The training set has $50,000$ images and the test set has $10,000$ images.
@@
ImageNet-100 is composed by randomly sampled 100 categories from ImageNet-1K (Deng et al., 2009). This dataset contains the following classes: n01498041, n01514859, n01582220, n01608432, n01616318, n01687978, n01776313, n01806567, n01833805, n01882714, n01910747, n01944390, n01985128, n02007558, n02071294, n02085620, n02114855, n02123045, n02128385, n02129165, n02129604, n02165456, n02190166, n02219486, n02226429, n02279972, n02317335, n02326432, n02342885, n02363005, n02391049, n02395406, n02403003, n02422699, n02442845, n02444819, n02480855, n02510455, n02640242, n02672831, n02687172, n02701002, n02730930, n02769748, n02782093, n02787622, n02793495, n02799071, n02802426, n02814860, n02840245, n02906734, n02948072, n02980441, n02999410, n03014705, n03028079, n03032252, n03125729, n03160309, n03179701, n03220513, n03249569, n03291819, n03384352, n03388043, n03450230, n03481172, n03594734, n03594945, n03627232, n03642806, n03649909, n03661043, n03676483, n03724870, n03733281, n03759954, n03761084, n03773504, n03804744, n03916031, n03938244, n04004767, n04026417, n04090263, n04133789, n04153751, n04296562, n04330267, n04371774, n04404412, n04465501, n04485082, n04507155, n04536866, n04579432, n04606251, n07714990, n07745940.
@@
MNIST-C (Mu & Gilmer, 2019) is a corrupted version of MNIST data with different corruption types for benchmarking out-of-distribution robustness in computer vision.
@@
CIFAR-10-C is algorithmically generated, following the previous leterature (Hendrycks & Dietterich, 2018), from different corruptions for CIFAR-10 data including gaussian noise, defocus blur, glass blur, impulse noise, shot noise, snow, and zoom blur.
@@
ImageNet-100-C is algorithmically generated with Gaussian noise based on (Hendrycks & Dietterich, 2018) for the ImageNet-100 dataset (Deng et al., 2009).
@@
FashionMNIST (Xiao et al., 2017) consists of $70,000$ fashion products images from 10 categories, with $7,000$ images per category. There are $60,000$ training images and $10,000$ test images. The 10 categories include T-Shirt, Trouser, Pullover, Dress, Coat, Sandals, Shirt, Sneaker, Bad, and Ankle boots.
@@
SVHN (Netzer et al., 2011) is a real-world image dataset obtained from house numbers in Google Street View images. This dataset $73,257$ samples for training, and $26,032$ samples for testing with 10 classes.
@@
Places365 (Zhou et al., 2017) contains scene photographs and diverse types of environments encountered in the world. The scene semantic categories consist of three macro-classes: Indoor, Nature, and Urban.
@@
LSUN-C (Yu et al., 2015) and LSUN-R (Yu et al., 2015) are large-scale image datasets that are annotated using deep learning with humans in the loop. LSUN-C is a cropped version of LSUN and LSUN-R is a resized version of the LSUN dataset, which has no overlap categories with the CIFAR dataset (Krizhevsky et al., 2009).
@@
Textures (Cimpoi et al., 2014) refers to the Describable Textures Dataset, which contains a large dataset of visual attributes including patterns and textures. The subset we used has no overlap categories with the CIFAR dataset (Krizhevsky et al., 2009).
@@
iNaturalist (Van Horn et al., 2018) is a challenging real-world dataset with iNaturalist species, captured in a wide variety of situations. It has 13 super-categories and 5,089 sub-categories. We use the subset from (Huang & Li, 2021) that contains 110 plant classes that no category overlaps with IMAGENET-1K (Deng et al., 2009).
@@
OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$ OOD Acc.normal-$\uparrow$ ID Acc.normal-$\uparrow$ FPRnormal-$\downarrow$ AUROCnormal-$\uparrow$

### v2

@@ 4.3 Qualitative Insights
Figure 3 shows t-SNE visualizations (van2008visualizing) of the penultimate-layer feature embedding, for WOODS (c) vs. our method (d). The embeddings are extracted from the test split. The blue points denote the test ID set (CIFAR-10), the green points are the test samples from CIFAR-10-C, and the gray points are from SVHN. This visualization suggests that embeddings of CIFAR and CIFAR-C become more aligned with our margin-based optimization, which arguably leads to improved OOD generalization performance. This observation corroborates our theoretical insight in Section 3.3.
@@ 4.4 Experiments on ImageNet
In this section, we provide additional large-scale results on the ImageNet benchmark. We use ImageNet-100 as the in-distribution data ($\mathbb{P}_{\text{in}}$), with labels provided in Appendix C. For the covariate-shifted OOD data ($\mathbb{P}_{\text{out}}^{\text{covariate}}$), we use ImageNet-100-C with Gaussian noise in the experiment. For the semantic-shifted OOD data, we use the high-resolution natural images from iNaturalist (van2018inaturalist), with the same subset as MOS (huang2021mos). The wild data $\mathbb{P}_{\text{in}}$ is a mixture of ID data, covariate-shifted data ($\pi_{c}=0.5$), and semantic shifted data ($\pi_{s}=0.1$). We fine-tune ResNet-34 (he2016deep) (pre-trained on ImageNet) for 100 epochs, with an initial learning rate of 0.01 and a batch size of 64. Results in Table 5 suggest that our method can improve both ID and OOD accuracy compared to WOODS (the most competitive baseline).
@@ 5 Related Works
is of vital importance for machine learning models deployed in the open world. Recent advances in OOD detection can be broadly categorized into post hoc and regularization-based methods. In particular, post hoc methods (hendrycks2016baseline; liang2018enhancing; lee2018simple; liu2020energy; huang2021importance; sun2021react; sun2022out) focus on deriving test-time OOD scoring functions for a pre-trained classifier. Our proposed work is closer to another line of work (bevandic2018discriminative; hendrycks2018deep; malinin2018predictive; liu2020energy; du2021vos; ming2022posterior), which addresses the OOD detection problem by training-time regularization. For example, models are encouraged to give predictions with lower confidence (hendrycks2018deep) or higher energies (liu2020energy). These methods require access to a clean OOD dataset for training, which can be restrictive. To circumvent this, a recent work WOODS (katz2022training) first explored using wild mixture data consisting of the ID and semantically shifted OOD data $\mathbb{P}_{\text{wild}}:=(1-\pi)\mathbb{P}_{\text{in}}+\pi\mathbb{P}_{\text{out}}^{\text{semantic}}$. The crucial difference between our work and WOODS is whether the wild mixture data contains covariate-shifted data, which introduces new challenges not considered in prior work. As we show in this work, our formulation uniquely enables both OOD generalization and OOD detection, in one coherent framework.
@@
is a fundamental problem in machine learning, which aims to generalize to covariate-shifted data without any sample from the target domain (muandet2013domain; arjovsky2019invariant; bahng2020learning; wang2022generalizing; xie2020n). OOD generalization is more challenging compared to the classic domain adaptation problem (daume2006domain; blitzer2008learning; ben2010theory; ganin2015unsupervised; tzeng2017adversarial; redko2019advances; kang2019contrastive; kumar2020understanding; wang2022embracing), which assumes access to labeled samples from the target domain.
To highlight a few works, IRM (arjovsky2019invariant) and its variants (krueger2021out; ahuja2020invariant) aim to find invariant representation from different training environments via an invariant risk regularizer. GroupDRO (sagawa2019distributionally) and Probabilistic Group DRO (ghosal2023pdgro) minimize the worst-case training loss over a set of groups. zhou2020learning propose generating outlier samples of a novel domain, which are then used for improving the generalization of the classifier. Besides algorithm innovation, benchmark efforts have also been pursued by DomainBed (gulrajani2020search) and OoD-Bench (ye2022ood), which facilitates evaluation on OOD generalization. OS-SDG (zhucrossmatch) relies on one source domain with labels to train the model, while our framework can exploit unlabeled wild data naturally arising in the wild, which is a mixed composition of three data distributions. Different from previous literature, we focus on improving OoD robustness in classifiers by learning from the wild mixture data and building an OOD detector at the same time. To the best of our knowledge, we are the first work that leverages wild data for both OOD generalization and OOD detection purposes. Our framework also uniquely allows leveraging covariate-shifted data freely arising in the wild, without requiring any labeling.
@@
aims to leverage labeled data from a related domain (source domain) and improve the model performance for the target domain, where there exists category gap for label sets between the source and target domains (you2019universal). Several works have been proposed to address this problem (saito2020universal; fu2020learning; li2021domain; chen2022geometric; kundu2022subsidiary; chang2022unified; garg2022domain). UAN (you2019universal) presents a universal adaptation network that exploits both the domain similarity and prediction uncertainty of each sample for promoting common-class adaptation. DANCE (saito2020universal) proposes a domain adaptative neighborhood clustering technique for category shift-agnostic adaptation via entropy optimization. The work in chang2022unified proposes a unified optimal transport-based framework to encourage both global cluster discrimination and local consistency of samples. Different from prior works, we leverage both labeled in-distribution data and unlabeled wild data when training our model. Such unlabeled wild data naturally arise in real-world environments and have not been considered in prior literature.
@@
is a classic machine learning problem, which aims to learn classifiers from positive and unlabeled data (letouzey2000learning). Multiple prior works have been proposed for discussing PU learning (hsieh2015pu; zhao2022dist; acharya2022positive; chapel2020partial; xu2021positive). The work in (niu2016theoretical) proposes a theoretical comparison of positive-unlabeled learning against positive-negative learning based on the upper bounds of estimation errors.  du2015convex presents a convex formulation for PU learning by using different loss functions for positive and unlabeled samples. Margin-based PU learning (gong2018margin) introduces a provable positive margin-based PU learning algorithm for classification under the truncated linear distributions. There are two key differences between ours and PU learning: (1) PU learning only considers the task of distinguishing $\mathbb{P}_{\text{out}}$ (anomalous) and $\mathbb{P}_{\text{in}}$ (normal), not the task of doing classification simultaneously. We consider OOD detection which additionally requires learning a classifier for the distribution $\mathbb{P}_{\mathcal{X}\mathcal{Y}}$. (2) PU learning does not consider the generalization aspect under covariate-shifted OOD data, whereas our framework handles it in addition to semantic-shifted OOD data.
@@ Appendix B Experimental Details
[…]
@@
We use Wide ResNet (zagoruyko2016wide) with 40 layers and widen factor of 2. For evaluating the post hoc OOD detection baselines, we use the model trained with the CE loss on $\mathbb{P}_{\text{in}}$. We employ the same pre-trained model from the github: https://github.com/wetliu/energy_ood, which was trained on the complete CIFAR-10 dataset. Specifically, the model is trained using cross-entropy loss for 200 epochs. The learning rate is started at 0.1 and then decays by multiplier 0.1 at 100, 150, and 175 epochs. To facilitate easy comparison, the pre-trained model and baseline results are consistent with (katz2022training) (courtesy of Table 4).
@@
For OOD generalization baselines, we use the same network architecture, Wide ResNet-40-2 (zagoruyko2016wide), and train from scratch using respective losses. The baselines in Table 3 are trained on CIFAR-10 (krizhevsky2009learning) $50,000$ labeled training examples. We follow the default hyperparameter setting as in original papers whenever applicable. For Mixup (zhang2018mixup), we follow the original paper and set hyperparameter $\alpha=1$ to control the strength of interpolation between feature-target pairs. The $\lambda$ for IRM (arjovsky2019invariant) baseline is set to 100 and the $\lambda$ for VREx (krueger2021out) is set to 10 for the penalty weights. Following the original training configuration in WRN (zagoruyko2016wide), all the baselines are trained for 200 epochs with batch size 128. We use the SGD optimizer with an initial learning rate of 0.1. The learning rate decays by a factor of 10 after 60, 120, and 180 epochs. Weight decay is set to $10^{-4}$. All models are implemented in PyTorch 1.8.1. We evaluate the trained model on the CIFAR-10 test set (ID accuracy) and CIFAR-10-C (OOD accuracy).
@@ Appendix C Details of Datasets
MNIST (lecun1998mnist) is a large database of handwritten digits with 10 categories and is widely used in the field of machine learning. The MNIST contains $60,000$ training images and $10,000$ test images.
@@
CIFAR-10 (krizhevsky2009learning) contains $60,000$ color images with 10 classes. The training set has $50,000$ images and the test set has $10,000$ images.
@@
ImageNet-100 is composed by randomly sampled 100 categories from ImageNet-1K (deng2009imagenet). This dataset contains the following classes: n01498041, n01514859, n01582220, n01608432, n01616318, n01687978, n01776313, n01806567, n01833805, n01882714, n01910747, n01944390, n01985128, n02007558, n02071294, n02085620, n02114855, n02123045, n02128385, n02129165, n02129604, n02165456, n02190166, n02219486, n02226429, n02279972, n02317335, n02326432, n02342885, n02363005, n02391049, n02395406, n02403003, n02422699, n02442845, n02444819, n02480855, n02510455, n02640242, n02672831, n02687172, n02701002, n02730930, n02769748, n02782093, n02787622, n02793495, n02799071, n02802426, n02814860, n02840245, n02906734, n02948072, n02980441, n02999410, n03014705, n03028079, n03032252, n03125729, n03160309, n03179701, n03220513, n03249569, n03291819, n03384352, n03388043, n03450230, n03481172, n03594734, n03594945, n03627232, n03642806, n03649909, n03661043, n03676483, n03724870, n03733281, n03759954, n03761084, n03773504, n03804744, n03916031, n03938244, n04004767, n04026417, n04090263, n04133789, n04153751, n04296562, n04330267, n04371774, n04404412, n04465501, n04485082, n04507155, n04536866, n04579432, n04606251, n07714990, n07745940.
@@
MNIST-C (mu2019mnist) is a corrupted version of MNIST data with different corruption types for benchmarking out-of-distribution robustness in computer vision.
@@
CIFAR-10-C is algorithmically generated, following the previous leterature (hendrycks2018benchmarking), from different corruptions for CIFAR-10 data including gaussian noise, defocus blur, glass blur, impulse noise, shot noise, snow, and zoom blur.
@@
ImageNet-100-C is algorithmically generated with Gaussian noise based on (hendrycks2018benchmarking) for the ImageNet-100 dataset (deng2009imagenet).
@@
FashionMNIST (xiao2017fashion) consists of $70,000$ fashion products images from 10 categories, with $7,000$ images per category. There are $60,000$ training images and $10,000$ test images. The 10 categories include T-Shirt, Trouser, Pullover, Dress, Coat, Sandals, Shirt, Sneaker, Bad, and Ankle boots.
@@
SVHN (netzer2011reading) is a real-world image dataset obtained from house numbers in Google Street View images. This dataset $73,257$ samples for training, and $26,032$ samples for testing with 10 classes.
@@
Places365 (zhou2017places) contains scene photographs and diverse types of environments encountered in the world. The scene semantic categories consist of three macro-classes: Indoor, Nature, and Urban.
@@
LSUN-C (yu2015lsun) and LSUN-R (yu2015lsun) are large-scale image datasets that are annotated using deep learning with humans in the loop. LSUN-C is a cropped version of LSUN and LSUN-R is a resized version of the LSUN dataset, which has no overlap categories with the CIFAR dataset (krizhevsky2009learning).
@@
Textures (cimpoi2014describing) refers to the Describable Textures Dataset, which contains a large dataset of visual attributes including patterns and textures. The subset we used has no overlap categories with the CIFAR dataset (krizhevsky2009learning).
@@
iNaturalist (van2018inaturalist) is a challenging real-world dataset with iNaturalist species, captured in a wide variety of situations. It has 13 super-categories and 5,089 sub-categories. We use the subset from (huang2021mos) that contains 110 plant classes that no category overlaps with IMAGENET-1K (deng2009imagenet).
@@
[…]